At what rate should the earth rotate so that the apparent g at the equator becomes zero? What will be the length of the day in this situation?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. The apparent ‘g’ at equator becomes zero. i.e. g’ = g – ω^2 R = 0 or g = ω^2R or ω = √g/R = √9.8/6400 * 10^3 = √1.5 * 10^-6 = 1.2 * 10^-3 rad/s. T = 2π/ω = 2 * 3.14/1.2 * 10^-3 = 1.5 * 10^-6 sec. = 1.41 hour